OpenStudy (anonymous):

What is the simplified form of the sixth root of x times the sixth root of x times the sixth root of x times the sixth root of x? x to the two-thirds power x to the three halves power 4 times the sixth root of x the sixth root of 4x

OpenStudy (jdoe0001):

\(\large \bf \sqrt[6]{x}\cdot \sqrt[6]{x}\cdot \sqrt[6]{x}\cdot \sqrt[6]{x}\quad ?\)

OpenStudy (anonymous):

yes